# Advancing the Understanding and Evaluation of AR-Generated Scenes: When Vision-Language Models Shine and Stumble

### v2

@@ Introduction
Augmented reality (AR) has transformed how we interact with the world, unlocking groundbreaking applications across education, entertainment, and healthcare. However, some AR applications can negatively impact user experience, either unintentionally or through deliberate malice. For example, AR content that does not blend seamlessly with the physical environment—due to issues like spatial misalignment [22] or stylistic mismatches [11]—can reduce immersion and undermine the intended objectives of the AR experience. Additionally, some AR content may mislead or confuse users [18]. These challenges not only compromise user experience but also raise safety and ethical concerns. Therefore, evaluating the quality of AR content has become a critical challenge to ensure usability and safety in AR applications.
While prior studies have investigated methods for assessing general image quality [12, 21], evaluating AR scenes introduces unique challenges that surpass those of traditional image analysis. Unlike conventional approaches that primarily rely on mathematical metrics (e.g., SSIM [9]) or, more recently, machine learning models (e.g., IQAGPT [4]), the quality of an AR scene is deeply intertwined with the user’s subjective experience. This is affected by factors such as content placement, lighting, rendering quality, and the seamless integration of virtual elements into the physical environment. Typically, these aspects are assessed through user studies [14], which gather feedback after users interact with the AR app. However, this approach has notable drawbacks: experiencing an AR scene is often time-consuming, and variations in real-world environments or AR content during interactions can skew the evaluation. Real-time AR scene quality estimation offers a promising alternative, enabling immediate and holistic assessments. Yet, collecting real-time user feedback may not always be practical, as it could disrupt the interaction and detract from the experience. This highlights the need for objective and real-time methods to automatically evaluate AR content.
@@
Recent developments in Vision Language Models (VLMs) present potential solutions to the challenges of AR experience evaluation. These models enable a holistic, human-like understanding of complex scenes [24], excelling in context-aware analysis by capturing relationships between objects and generating detailed descriptions that align with human perception. However, before using VLMs for AR content evaluation, it is essential to understand VLMs’ ability toward AR content understanding. Examples in Figure 1 present AR scenes paired with GPT’s responses to the question, “Is this an augmented reality (AR) scene? If yes, what virtual content does it include, and could you explain why each element is considered virtual?” These examples suggest that VLMs can recognize AR images but are sometimes unreliable. To address this, in this work, we provide an assessment of the feasibility of using modern VLMs for evaluating AR content. Toward this goal, we collected AR images from different sources and tested three state-of-the-art VLMs, GPT [16], Gemini [7], and Claude [1], with different prompt questions on the images. The contributions of this paper include:
  • We curate and release a dataset [^1] https://github.com/ARResearch-1/DiverseAR-Dataset, DiverseAR, comprising 318 images collected from two commercial AR platforms (Amazon and Scaniverse), three AR applications previously developed by our lab running on Magic Leap, Android, and HoloLens [17, 23, 6], and two AR applications specifically created for this project running on Apple Vision Pro and Android. This dataset is the first of its kind for assessing the capabilities of VLMs in identifying and describing AR content, encompassing diverse AR scenarios.
  • We test VLMs on identifying and describing AR content. Results show that VLMs are good at recognizing and describing obvious and typical AR content, reaching a perception TPR of up to 93% and description TPR of up to 71% across varying AR scene complexity levels within the DiverseAR dataset.
@@ 1 Related Work
Despite their promise, concerns persist regarding VLMs’ perception and reasoning capabilities. Studies have shown that VLMs can hallucinate, producing information not grounded in the data [20]. Additionally, VLMs exhibit limited ability to comprehend depth information [3], further raising questions about their reliability in evaluating the quality of AR experiences, particularly in aspects such as spatial alignment, contextual integration, and rendering quality. To address these issues, this work evaluates the ability of VLMs to analyze and interpret AR images, offering a pioneering investigation into their effectiveness in AR content evaluation.
@@ 2.1 Motivation and Collection Process
To evaluate the AR scene understanding capabilities of VLMs, we curated the DiverseAR dataset, specifically designed to capture a broad spectrum of AR scenarios. The dataset consists of 298 AR images collected from diverse sources and environments. It includes 23 images captured using a custom-developed Apple Vision Pro AR application in laboratory and kitchen environments, and 151 images collected from a custom-developed Android AR application in bedroom and dining room environments. Additionally, 42 images were created to explore AR-specific research topics, such as attention patterns[17], virtual content arrangements[23], and surgical guidance[6]. The dataset also features 7 images of glass objects obtained from the Amazon app’s AR view and 46 images collected from the Scaniverse[15] app’s AR view, captured in laboratory, kitchen, and dining room environments. Finally, 29 images were sourced from a website showcasing AR advertisement videos [5]. Additionally, we included 20 non-AR images to supplement the dataset.
@@ 2.2 Dataset Composition
The DiverseAR dataset encompasses a wide spectrum of AR and non-AR scenarios, showcasing diverse characteristics of both virtual and real-world content. Table 1 summarizes the key characteristics represented in the dataset, while Figure 2 illustrates representative examples of these characteristics. The dataset includes single or multiple instances of identical and varied real and virtual objects across various AR images. These objects span numerous classes, including toys, food, shoes, plants, laptops, containers, etc.
By offering this diversity in both content and context, the DiverseAR dataset enables comprehensive evaluation of VLMs in interpreting simple and complex AR scenes.
@@ 3 Methods
To evaluate the AR scene understanding capabilities of VLMs and identify factors influencing their performance, we assess three commercial VLMs using our DiverseAR dataset: 1) GPT-4o-2024-08-06 by OpenAI; 2) Claude-3.5-Sonnet-20241022 by Anthropic; 3) Gemini-1.5-Pro-002 by Google. These models represent state-of-the-art capabilities in vision-language understanding and generation. We also design targeted prompts to assess their perception of AR scenes and the accuracy of their descriptions. We define three levels of AR scene complexity and categorize AR samples based on VLM responses. These categories enable a quantitative analysis of VLM performance in recognizing and describing AR scenes.
@@ 3.1 Prompt Design for AR Scene Understanding
To evaluate the AR scene understanding abilities of VLMs, we design two types of prompts: general and task-aware. These prompts aim to assess the VLMs’ capacity to generate comprehensive textual descriptions of the entire image.
@@
This prompt is designed to assess the model’s ability to generate a general description of the image without any specific guidance toward identifying AR content. It tests whether the model can naturally recognize and articulate the presence of AR elements without explicit textual cues.
Task-aware Image Captioning Prompt: 1) Without mentioning AR:Does this image include virtual content superimposed on the real world? If yes, could you list all the virtual elements present and explain why each one is considered virtual?” 2) With mentioning AR:Is this an augmented reality (AR) scene? If yes, what virtual content does it include, and could you explain why each element is considered virtual?
These task-aware images captioning prompts are designed to assess the model’s ability to leverage guidance for more accurate identification and description of AR-specific elements. Since the results for the prompt mentioning AR and the prompt not mentioning AR were found to be similar in VLMs’ responses, we chose to use the task-aware prompt without mentioning AR in our experiments for consistency and simplicity.
@@ 3.2 Classification of AR Scene Complexity Levels
To evaluate VLM performance under varying levels of AR scene complexity, we define three AR scene difficulty levels as follows:
@@ 3.3 Categorization of VLM Responses
To analyze the factors that affect VLMs’ performance, we categorize AR samples into five groups based on the VLMs’ response:
@@ 3.4 Evaluation Metrics
where $N_{1}$, $N_{2}$, $N_{3}$, $N_{4}$, and $N_{5}$ denote the number of image samples in Category 1, 2, 3, 4, and 5, respectively. $TPR_{P}$ represents the proportion of AR images identified as containing AR elements among all AR images, regardless of the accuracy of the AR scene descriptions. Similarly, $TNR_{P}$ indicates the proportion of non-AR images described without referencing AR elements among all non-AR images, irrespective of the accuracy of the non-AR scene descriptions.
@@
$TPR_{D}$ represents the percentage of AR images that are both correctly identified as containing AR elements and accurately described in terms of the AR scene, out of all AR images. This metric, alongside perception metrics, allows for a detailed evaluation of the VLMs’ strengths and limitations in AR scene understanding. By focusing on the quality of both recognition and description, we aim to provide detailed insights into VLM performance across varying levels of AR scene complexity. These results are analyzed further in Section 4.
@@ 4 Results
Figure 3: The True Positive Rate for Perception ($TPR_{P}$) of the three VLMs across varying complexity levels using G (general) and T (task-aware) image captioning prompt. VLMs show a consistent decline in performance as AR scene complexity increases.
Figure 4: The True Positive Rate for Description ($TPR_{D}$) of the three VLMs across varying AR scene complexity levels using G (general) and T (task-aware) image captioning prompt. Notably, GPT outperforms both Gemini and Claude on medium and hard AR scenes using prompt T.
We present the evaluation results of VLMs’ AR scene understanding capabilities, focusing on their perception and description performance. An in-depth categorical analysis of VLM responses is also discussed. Furthermore, we compare the performance of VLMs with that of human counterparts.
@@ 4.1 Performance of VLMs’ AR Understanding Abilities
Perception Performance: We quantitatively evaluate the ability of VLMs to perceive AR scenes across varying levels of scene complexity and the entire DiverseAR dataset using the prompts outlined in  Section 3.
Across the dataset, the Perception True Negative Rate ($TNR_{P}$) for all three VLMs consistently remains at 100%. This demonstrates that VLMs rarely misclassify real-world images as AR content, showcasing their robustness in distinguishing non-AR scenes.
When comparing the Perception True Positive Rate ($TPR_{P}$) using the task-aware prompt (T) and the general prompt (G), we observe significant improvements with prompt T across all models, as illustrated in Figure 3. Performance improves by a range of 30.5% to 46.7%, highlighting the importance of explicitly specifying AR-related tasks within prompts. Among the 3 models we evaluate, Gemini consistently achieves the highest $TPR_{P}$, with 61.4% under prompt G and 93.3% under prompt T, outperforming GPT and Claude. These results underscore Gemini’s superior performance and its potential suitability for AR scene recognition tasks.
@@
The prompt T significantly improves description accuracy compared to the prompt G, as shown in Figure 4. For instance, GPT’s Description True Positive Rate ($TPR_{D}$) increases from 38.9% with prompt G to 73.8% with prompt T, emphasizing the importance of tailored prompts. While Gemini outperforms GPT and Claude under prompt G, GPT achieves the highest $TPR_{D}$ under prompt T (73.8%), exceeding Gemini’s performance (71.8%) by 2%. This result indicates that GPT has the potential to more effectively leverage explicit task prompts when analyzing AR scenes. In contrast, Claude consistently underperforms, with $TPR_{D}$ as low as 14.1% under prompt G, improving to only 34.9% under prompt T. Detailed analysis reveals that Claude often generates high-level summaries of objects in the scene without capturing the relationships among them, limiting its ability to accurately describe AR-specific content.
@@ 4.2 Categorical Analysis of VLMs’ Responses
We analyzed AR samples across the five response categories defined in Section 3, examining key observations, contributing factors, and possible reasons for the VLMs’ behavior.
Category 1: Accurate AR Recognition. In this category, VLMs demonstrate strong performance in identifying AR objects. We observe that AR content with characteristics such as text overlays, low-quality rendering (e.g., uniform pixel distributions), or implausible physical attributes (e.g., lighting, shape, or size mismatches, unnatural object placement like floating objects, or unrealistic transparency) is often recognized correctly, as shown in Figure 2(k, l, e, d). These attributes are prevalent in common AR scenarios and are likely well-represented in the VLMs’ training datasets, making the models adept at detecting them.
Interestingly, certain visual cues enhance recognition accuracy. For instance, the presence of QR codes or hands provides strong contextual clues. As shown in Figure 5, a virtual sign with a QR code is correctly identified, while the same virtual sign without it is misclassified. Similarly, as seen in Figure 6, when hands are included in the scene, VLMs are more likely to recognize the AR content correctly, whereas their absence leads to recognizing the real grid as virtual content. Additionally, implausible physical interactions, such as a pot placed on a cat, further help the models identify AR content. The ability of VLMs to identify these features can be attributed to the self-attention mechanism in transformer-based architectures [8], which processes an image as a sequence of patches, capturing global context and spatial relationships. This design enables the models to effectively understand and contextualize relationships, actions, and attributes within an image.
Figure 5: These images feature an AR-generated “Exit” sign integrated into real-world contexts. With the prompt T, GPT accurately identifies the sign with a QR code (bottom) but misclassifies it as a real object without it (top).
@@
Category 2: Partial AR Recognition. In this category, VLMs misinterpret AR content, in three primary ways. First, real objects are sometimes recognized as virtual. For instance, as shown in Figure 7, a real pair of sneakers is identified correctly when no virtual sneakers are present. However, adding virtual sneakers to the scene causes the VLM to misclassify the real sneakers as virtual. Second, virtual content is often only partially recognized. For example, in mixed scenes, a virtual toy elephant and a virtual pot are not recognized as virtual objects, as illustrated in Figure 8. Additionally, shadows rendered on virtual objects can enhance their realism, allowing them to evade detection. Lastly, virtual objects may be misclassified entirely or have their properties misunderstood; for example, a toy butterfly might be incorrectly identified as another type of object, possibly due to gaps in the training data. These issues are most common when virtual and real elements are combined, creating ambiguity that misleads VLMs. The model’s response to a single object can vary depending on its surrounding elements, suggesting that these errors may stem from the spatial reasoning capabilities of transformer-based models.
@@
Figure 8: This image features an AR-generated pot, heater, apple, toy elephant, and toy android integrated into real-world contexts. With prompt T, GPT fails to identify a virtual toy elephant and a virtual pot with realistic shadows as virtual objects.
Category 3: Missed AR Recognition. Failures in this category arise when realistic virtual content adheres to physical laws or is seamlessly integrated into real-world settings. These scenarios often mislead VLMs, making it difficult for them to identify virtual elements. We observe that virtual objects with highly realistic shadows often evade detection, particularly when placed in environments lacking strong visual cues, such as dark surfaces, low-light settings, or inverse lighting directions, as shown in Figure 2(j). Similarly, virtual objects aligned with real-world edges or positioned on digital screens can further mislead VLMs, as seen in Figure 9(a). Additionally, we notice that virtual objects with inadequate transparency are often misidentified as real, as shown in Figure 9(b). Some virtual content with low rendering quality may be misclassified as realistic “fake objects”, as such “fake objects” are often encountered in daily life, as seen in Figure 9(c). Moreover, objects that defy physical laws—such as floating slightly but remaining attached to the surfaces of real objects—still appear plausible enough to evade detection, as illustrated in Figure 2(j). Virtual objects that do not typically appear in AR settings further exacerbate the recognition problem, as shown in Figure 2(h).
@@
These issues highlight several limitations of VLMs. For instance, they struggle with depth perception in 2D images and have trouble interpreting situations that challenge common-sense knowledge [25], making it difficult to identify violations of some physical laws. Additionally, we hypothesize that many models may have limited exposure to diverse or atypical AR-specific samples during training, potentially hindering their ability to effectively understand AR scenes.
Category 4: False AR Detection. No instances of false AR detection were observed, showcasing the model’s robustness in accurately distinguishing non-AR images. To further investigate the generalizability of this observation, our ongoing work involves creatively designing real-world images with implausible lighting and placements. Examples include shining a light onto a real apple or attaching a pen unnaturally to a shelf.

### v3

@@ Introduction
Augmented Reality (AR) is poised to transform how we interact with the world, unlocking groundbreaking applications across education, entertainment, and healthcare. However, some AR applications can negatively impact user experience, either unintentionally or through deliberate malice. For example, AR content that does not blend seamlessly with the physical environment can reduce immersion and undermine the intended objectives of the AR experience, due to issues like spatial misalignment [22] or stylistic mismatches [11]. Additionally, some AR content may mislead or confuse users [18]. These challenges not only compromise user experience but also raise safety and ethical concerns. Therefore, evaluating the quality of AR content has become a critical challenge to ensure usability and safety in AR applications.
While prior studies have investigated methods for assessing general image quality [12, 21], evaluating AR scenes introduces unique challenges that surpass those of traditional image analysis. Unlike conventional approaches that primarily rely on mathematical metrics (e.g., SSIM [9]) or, more recently, machine learning models (e.g., IQAGPT [4]), the quality of an AR scene is deeply intertwined with the user’s subjective experience. This is affected by factors such as content placement, lighting, rendering quality, and the seamless integration of virtual elements into the physical environment. Typically, these aspects are assessed through user studies [14], which gather feedback after users interact with the AR app. However, this approach has some drawbacks: experiencing an AR scene is often time-consuming, and variations in real-world environments or AR content during interactions can skew the evaluation. Real-time AR scene quality estimation offers a promising alternative, enabling immediate and holistic assessments. Yet, collecting real-time user feedback may not always be practical, as it could disrupt the interaction and detract from the experience. This highlights the need for objective and real-time methods to automatically evaluate AR content.
@@
Recent developments in Vision Language Models (VLMs) present potential solutions to the challenges of AR experience evaluation. These models enable a holistic, human-like understanding of complex scenes [24], excelling in context-aware analysis by capturing relationships between objects and generating detailed descriptions that align with human perception. However, it is essential to understand VLMs’ ability toward AR content understanding before using them for AR content evaluation. Examples in Figure 1 present AR scenes together with GPT’s responses to the question, “Is this an augmented reality (AR) scene? If yes, what virtual content does it include, and could you explain why each element is considered virtual?” These examples suggest that VLMs can recognize AR images but are sometimes unreliable. To address this, in this work, we provide an assessment of the feasibility of using modern VLMs for evaluating AR content. Toward this goal, we collected AR images from different sources and tested three state-of-the-art VLMs, GPT [16], Gemini [7], and Claude [1], with different prompt questions on these images. The contributions of this paper include:
  • We curate and release a dataset [^1] https://github.com/ARResearch-1/DiverseAR-Dataset, DiverseAR, comprising 318 images collected from one public website [5], two commercial AR platforms (Amazon and Scaniverse), three AR applications previously developed by our lab running on Magic Leap, Android, and HoloLens [17, 23, 6], along with two custom-built AR applications for this project, running on Apple Vision Pro and Android. To the best of our knowledge, this dataset is the first of its kind for assessing the capabilities of VLMs in identifying and describing AR content, encompassing diverse AR scenarios.
  • We test VLMs on identifying and describing AR content. Results show that VLMs are effective in recognizing and describing obvious and typical AR content, reaching a perception TPR of up to 93% and description TPR of up to 71% across varying AR scene complexity levels within the DiverseAR dataset.
@@ 1 Related Work
Despite their promise, concerns persist regarding VLMs’ perception and reasoning capabilities. Studies have shown that VLMs can hallucinate, producing information not grounded in the data [20]. Additionally, VLMs exhibit limited ability to comprehend depth information [3], further raising questions about their reliability in evaluating the quality of AR experiences, particularly in aspects such as spatial alignment and contextual integration. In light of these issues, this work evaluates the ability of VLMs to analyze and interpret AR images, offering a pioneering investigation into their effectiveness in AR content evaluation.
@@ 2.1 Motivation and Collection Process
To evaluate the AR scene understanding capabilities of VLMs, we curated the DiverseAR dataset, specifically designed to capture a broad spectrum of AR scenarios. The dataset consists of 298 AR images collected from diverse sources and environments. It includes 23 images captured using a custom-developed Apple Vision Pro AR application in laboratory and kitchen environments, and 151 images collected from a custom-developed Android AR application in bedroom and dining room environments. Additionally, 42 images were collected during the exploration of AR-specific research topics, such as attention patterns[17], virtual content arrangements[23], and surgical guidance[6]. The dataset also features 7 images from the Amazon app’s AR view and 46 images collected from the Scaniverse[15] app’s AR view, captured in laboratory, kitchen, and dining room environments. Finally, 29 images were sourced from a website showcasing AR advertisement videos [5]. Additionally, we included 20 non-AR images to supplement the dataset.
@@ 2.2 Dataset Composition
The DiverseAR dataset encompasses a wide spectrum of AR and non-AR scenarios, showcasing diverse characteristics of both virtual and real-world content. Table 1 summarizes the key characteristics represented in the dataset, while Figure 2 illustrates representative examples of these characteristics. The dataset includes single or multiple instances of identical and varied real and virtual objects across various AR images. These objects span various classes, such as toys, food, shoes, plants, laptops and containers.
By offering this diversity in both content and context, the DiverseAR dataset enables an insightful evaluation of VLMs in interpreting simple and complex AR scenes.
@@ 3 Methods
To evaluate the AR scene understanding capabilities of VLMs and identify factors influencing their performance, we assess three commercial VLMs using our DiverseAR dataset: 1) GPT-4o-2024-08-06 by OpenAI; 2) Claude-3.5-Sonnet-20241022 by Anthropic; 3) Gemini-1.5-Pro-002 by Google. These models represent state-of-the-art capabilities in vision-language understanding and generation. We also design targeted prompts to assess the accuracy of their perception and description of AR scenes. Additionally, we define three levels of AR scene complexity based on AR content attributes and categorize AR samples into five groups based on VLM responses. These categories facilitate a detailed quantitative analysis of VLM performance in recognizing and describing AR scenes.
@@ 3.1 Prompt Design for AR Scene Understanding
To assess the capacity of VLMs to generate comprehensive textual descriptions of images, we design two types of prompts: general and task-aware.
@@
This prompt is designed to assess the model’s ability to generate a general caption of the image without any specific guidance toward identifying AR content. It tests whether the model can naturally recognize and describe AR elements without explicit textual cues.
Task-aware Image Captioning Prompts: 1) Without explicitly referencing AR:Does this image include virtual content superimposed on the real world? If yes, could you list all the virtual elements present and explain why each one is considered virtual?” 2) Explicitly referencing AR:Is this an augmented reality (AR) scene? If yes, what virtual content does it include, and could you explain why each element is considered virtual?
These task-aware image captioning prompts are designed to assess the model’s ability to leverage guidance for more accurate identification and description of AR-specific elements. Since the responses from VLMs were similar for both task-aware prompts, we chose to use the prompt without explicitly referencing AR in our experiments for consistency and simplicity.
@@ 3.2 Classification of AR Scene Complexity Levels
To analyze VLM performance across varying complexities of AR scenes, we define three distinct levels based on the attributes of virtual elements as follows:
@@ 3.3 Categorization of VLM Responses
To facilitate the analysis of factors affecting VLM performance, we categorize AR samples into five groups based on the VLMs’ responses:
@@ 3.4 Evaluation Metrics
where $N_{1}$, $N_{2}$, $N_{3}$, $N_{4}$, and $N_{5}$ denote the number of images in Category 1, 2, 3, 4, and 5, respectively. $TPR_{P}$ represents the proportion of AR images identified as containing AR elements among all AR images, regardless of the accuracy of the AR scene descriptions. Similarly, $TNR_{P}$ indicates the proportion of non-AR images described without referencing AR elements among all non-AR images, irrespective of the accuracy of the non-AR scene descriptions.
@@
$TPR_{D}$ represents the percentage of AR images that are both correctly identified as containing AR elements and accurately described in terms of the AR scene, out of all AR images. This metric, alongside perception metrics, allows for a detailed evaluation of the VLMs’ strengths and limitations in AR scene understanding. By focusing on the quality of both perception and description, we aim to provide detailed insights into VLM performance across varying levels of AR scene complexity. These results are analyzed further in Section 4.
@@ 4 Results
Figure 3: The True Positive Rate for Perception ($TPR_{P}$) of the three VLMs across varying complexity levels using G (general) and T (task-aware) image captioning prompts. VLMs show a consistent decline in performance as AR scene complexity increases.
Figure 4: The True Positive Rate for Description ($TPR_{D}$) of the three VLMs across varying AR scene complexity levels using G (general) and T (task-aware) image captioning prompts. Notably, GPT outperforms both Gemini and Claude on medium and hard AR scenes using prompt T.
We present the results of VLMs’ AR scene understanding capabilities, focusing on their perception and description performance. We also provide an in-depth categorical analysis of VLM responses and compare their performance with that of human counterparts.
@@ 4.1 Performance of VLMs’ AR Understanding Abilities
Perception Performance:
We evaluate VLMs’ ability to perceive AR scenes at different complexity levels, using the DiverseAR dataset as well as the prompts defined in Section 3.
Across the dataset, the Perception True Negative Rate ($TNR_{P}$) remains consistently at 100% for all models. This demonstrates that VLMs rarely misclassify real-world images as AR content, showcasing their robustness in distinguishing non-AR scenes.
When comparing the Perception True Positive Rate ($TPR_{P}$) using the task-aware prompt (T) and the general prompt (G), we observe significant improvements with prompt T across all models, as illustrated in Figure 3. We observe performance improvements ranging from 30.5% to 46.7%, highlighting the importance of explicitly specifying AR-related textual cues. Among the three models we evaluated, Gemini consistently achieves the highest $TPR_{P}$, with 61.4% under prompt G and 93.3% under prompt T, outperforming GPT and Claude. These results underscore Gemini’s superior performance and its potential suitability for AR scene perception tasks.
@@
The prompt T significantly improves description accuracy compared to the prompt G, as shown in Figure 4. For instance, GPT’s Description True Positive Rate ($TPR_{D}$) increases from 38.9% with prompt G to 73.8% with prompt T, emphasizing the importance of tailored prompts. While Gemini outperforms GPT and Claude under prompt G, GPT achieves the highest $TPR_{D}$ under prompt T (73.8%), surpassing Gemini (71.8%) by 2%. This result indicates that GPT has the potential to leverage explicit task prompts more effectively when analyzing AR scenes. In contrast, Claude consistently underperforms, with $TPR_{D}$ as low as 14.1% under prompt G, improving to only 34.9% under prompt T. Detailed analysis reveals that Claude often generates high-level summaries of objects in the scene without capturing the relationships among them, limiting its ability to accurately describe AR-specific content.
@@ 4.2 Categorical Analysis of VLMs’ Responses
We analyze AR samples across the five response categories defined in Section 3, examining key observations, contributing factors, and potential reasons for the VLMs’ behavior.
Category 1: Accurate AR Recognition. In this category, VLMs demonstrate strong performance in identifying AR objects. We observe that AR content with characteristics such as text overlays, low-quality rendering (e.g., uniform pixel distributions), or implausible physical attributes (e.g., lighting, shape, size mismatches, unnatural object placement like floating objects, or unrealistic transparency) is often recognized correctly, as shown in Figure 2(k, l, e, d). These attributes are prevalent in common AR scenarios and are likely well-represented in the VLMs’ training datasets, making the models adept at detecting them.
Interestingly, certain visual cues enhance recognition accuracy. For instance, the presence of QR codes or hands provides strong contextual clues. As shown in Figure 5, a virtual sign with a QR code is correctly identified, while the same virtual sign without it is misclassified. Similarly, Figure 6 shows that the presence of hands improves AR recognition, while their absence causes VLMs to misclassify real grids as virtual content. Additionally, implausible physical interactions, such as a pot placed on a cat, further help the models identify AR content. The ability of VLMs to identify these features can be attributed to the self-attention mechanism in transformer-based architectures [8], which processes an image as a sequence of patches, capturing global context and spatial relationships. This design enables the models to effectively understand and contextualize relationships, actions, and attributes within an image.
Figure 5: These images feature an AR-generated “Exit” sign integrated into real-world contexts. With the prompt T, GPT accurately identifies the sign with a QR code (bottom) but misclassifies it as a real object without the code (top).
@@
Category 2: Partial AR Recognition. In this category, VLMs misinterpret AR content, in three primary ways. First, real objects are recognized as virtual. For instance, as shown in Figure 7, a real pair of sneakers is identified correctly when no virtual sneakers are present. However, adding virtual sneakers to the scene causes the VLM to misclassify the real sneakers as virtual. Second, virtual content is only partially recognized. For example, in a mixed scene, a virtual toy elephant and a virtual pot are not recognized as virtual objects, as illustrated in Figure 8. Shadows rendered on these virtual objects can enhance their realism, allowing them to evade detection. Lastly, virtual objects may be misclassified entirely or have their properties misunderstood. For example, a toy butterfly might be incorrectly identified as another type of object, possibly due to gaps in the training data. These issues are most common when virtual and real elements are combined, creating ambiguity that misleads VLMs. Consequently, the model’s response to a certain object can vary depending on its surrounding elements, suggesting that these errors likely stem from the spatial reasoning capabilities of transformer-based models.
@@
Figure 8: This image features an AR-generated pot, heater, apple, toy elephant, and toy android integrated into a real-world context. With prompt T, GPT fails to identify a virtual toy elephant and a virtual pot with realistic shadows as virtual objects.
Category 3: Missed AR Recognition. Failures in this category arise when realistic virtual content adheres to physical laws or is seamlessly integrated into real-world settings. These scenarios can mislead VLMs, making it difficult for them to identify virtual elements. For instance, virtual objects with realistic shadows frequently evade detection, especially in environments with weak visual cues. These include dark surfaces, low-light conditions, or inverse lighting directions (Figure 2(j)). Similarly, virtual objects positioned on digital screens can also mislead VLMs, as seen in Figure 9(a).
We also observe several interesting cases within this category. For example, virtual objects with inadequate transparency are misidentified as real (Figure 9(b)). Some virtual content with low rendering quality may be misclassified as realistic “fake objects”, as such “fake objects” are often encountered in daily life (Figure 9(c)). Likewise, objects that defy physical laws—such as floating slightly but remaining attached to the surfaces of real objects—still appear plausible enough to evade detection (Figure 2(j)). Moreover, virtual objects that do not typically appear in AR settings further exacerbate the recognition problem (Figure 2(h)).
@@
These observations highlight several limitations of VLMs. For instance, they struggle with depth perception in 2D images and have trouble interpreting situations that challenge common-sense knowledge [25], making it difficult to identify violations of some physical laws. Additionally, we speculate that many models may have limited exposure to diverse or atypical AR-specific samples during training, potentially hindering their ability to effectively understand AR scenes.
Category 4: False AR Detection. No instances of false AR detection were observed, showcasing the model’s robustness in accurately distinguishing non-AR images. To further investigate the generalizability of this observation, our ongoing work involves constructing real-world scenarios with implausible lighting and object placements. Examples include shining a light onto a real apple or attaching a pen unnaturally to a shelf.